vector u= 2i-j vector v= -2i+3J-3K find the component vector u perpendicular to v
Answer #1

2 +3435 Given vectors are U = 2i-j, V=-2i+3; -3k Component vector u perpendiculas tu v = U-XXV (j) (-24 3 - 3k) 2 -j - (-2i +